The temperature of the sun is doubled,the rate of energy received on earth will be increased by a factor of

Grade:12th Pass

3 Answers

shiddhant bhattacharya
11 years ago

16

Vikas TU
11 years ago

Ei = 3/2 x k x (T)................(a)

Ef = 3/2 x k x (2T).................(b)

 

(a)/(b) = Ei/Ef = 1/2

                2Ei = Ef

     ANS. factor of 2.

 

 

PLEASE  APPROVE!          

Pratibha Rani
10 years ago

sixteen